Find the MacLaurin series for the function f(x) = 12, ο ΣΩ(+ 1): ο ο Σ(-1) 0η Σε ο 2) η -0
Answer #1

given fox)= I - i that maclaurin we know given by series is f(x) = f(0) + flox+ flox? + f(0)33. ------ -ci Now I-X fa) = (1-0put the above values in eq is) IN f(x) = 1 + lxx + 2x126x21 2137 +244- 42 fx) = 1 + 2 + x2 + x3 + x4 +----- so we con corite